# Cascading and Proxy Membership Inference Attacks

### v2

@@ Abstract
publicationid: pubid: Network and Distributed System Security (NDSS) Symposium 2026 23-27 February 2026, San Diego, CA, USA ISBN 979-8-9894372-8-3 https://dx.doi.org/10.14722/ndss.2026.250661 www.ndss-symposium.org
@@ V-A Experimental Setup
  • LiRA [9] exploits behavioral discrepancies in the query instance by modeling its loss distribution as a Gaussian estimate and uses a likelihood ratio test to compute the membership score.
@@ V-E Additional Investigations
Attack on Non-image Datasets. We also evaluate the attack performance on two non-image datasets to demonstrate the generality of the proposed attacks. Specifically, we train multilayer perceptrons (MLPs) on the Purchase and Texas datasets [1]. For $\mathsf{CMIA}$, we use LiRA as the base model and apply the cascading framework. For $\mathsf{PMIA}$, we employ the Wasserstein distance to select the top 10 proxy models for each query instance. The attack performance under both adaptive and non-adaptive settings is presented in Table VII. As shown, both $\mathsf{CMIA}$ and $\mathsf{PMIA}$ consistently outperform existing attacks, particularly in TPR at low FPR. It is worth mentioning that the performance of all evaluated MIAs is notably lower on these non-image datasets compared to image datasets such as CIFAR-100, despite both having 100 classes. We attribute this to the smaller generalization gap in models trained on non-image datasets. For instance, as shown in Table XV, the Purchase model achieves training and validation accuracies of 0.96 and 0.80, respectively. In contrast, the CIFAR-100 model shows a much larger gap, with a training accuracy of 0.92 and a validation accuracy of only 0.66. Since MIAs typically exploit overfitting for attack [9], they are less effective against better-generalized models, as observed in the non-image datasets.
@@
Attack Against DP-SGD. Machine learning with differential privacy [43] is an effective defense mechanism against privacy attacks, including MIAs. We follow [9, 8] and assess the effectiveness of the DP-SGD against our attack. Specifically, we fix the clipping norm $C$ as 10 and test $\mathsf{CMIA}$ and $\mathsf{PMIA}$ on a ResNet50 model trained on CIFAR-10. As shown in Table X, even when only the gradient norm is applied without adding noise, the model’s accuracy and the effectiveness of our attack are significantly reduced. We observe that as the noise level increases, the improvement of $\mathsf{CMIA}$ compared to the base attack diminishes. This is because $\mathsf{CMIA}$ relies on identifying highly probable samples to perform cascading attacks on remaining instances. When selecting reliable anchors becomes difficult, the benefit of the cascading framework decreases. Nevertheless, $\mathsf{CMIA}$ consistently outperforms the base attack under all evaluated DP settings. We also evaluate the impact of DP-SGD in the non-adaptive setting in Table XI and observe similar patterns: as the privacy level increases, attack performance drops for all methods, and the gap between attacks narrows. Nevertheless, $\mathsf{PMIA}$ still outperforms existing MIAs in the low false-positive regime.
@@ VI Related Work
Non-Adaptive MIAs. Some studies [14, 12] argue that adaptive attacks are inefficient due to the need to train shadow models for every batch of queries. In response, a series of non-adaptive MIAs have been proposed to relax this assumption by training shadow models (or forgoing shadow training entirely) before accessing the query set. Several techniques have been introduced for non-adaptive MIAs, such as score functions [15, 39], difficulty calibration [24, 37, 8], loss trajectory [13, 56], quantile regression [12], and hypothesis testing [38, 14]. However, these attacks deviate from the membership posterior odds ratio test, leading to suboptimal attack performance.
Recent research [9, 57] has emphasized evaluating attacks by calculating their true positive rate at a significantly low false positive rate, and most existing MIAs perform poorly under this evaluation paradigm. In Section IV-A we provide theoretical insights into the rationale behind these evaluation metrics. While some literature [58] explores instance vulnerabilities to MIAs, it does not propose an attack that leverages them. In contrast, we theoretically analyze membership dependencies and introduce a framework that exploits these vulnerabilities and dependencies to mount a more powerful attack. Moreover, the dependencies we investigate differ from group-level MIAs [59, 60], which consider the inference target as a set of instances. Instead, our focus lies in exploring the membership dependencies between inference targets.
@@ VII Discussion & Conclusion
In this paper, we provide a new formulation of the MIA game and categorize existing MIAs into two categories: adaptive and non-adaptive. Guided by theoretical analyses of joint membership estimation and the posterior odds test, we propose two attacks (i.e., $\mathsf{CMIA}$ and $\mathsf{PMIA}$), one for each setting. Extensive experiments demonstrate the efficacy of the proposed attacks.
Our work has several limitations. First, the conditional shadow training strategy of $\mathsf{CMIA}$ limits its applicability for attacks that do not rely on shadow training. Second, our evaluations mainly focus on classification models, leaving the performance on generative models [61, 62, 63] unexamined. Nevertheless, our work provides both theoretical insights and practical approaches, offering new directions for MIAs.
@@
This paper focuses on membership inference attacks using publicly available datasets, such as MNIST and CIFAR-10, that do not contain personally identifiable information. We do not use any private or sensitive data, nor do we target specific individuals or proprietary models. Although our proposed attacks demonstrate improved attack performance, we show that established defenses like DP-SGD remain an effective mitigation strategy. We emphasize that the responsible use of such attack methods should be used responsibly to strengthen the security and privacy of ML systems, rather than maliciously exploiting them. We encourage the research community to use our findings in a manner that promotes ethical research and enhances privacy protections for users and data subjects.
@@ A-A Convergence of Joint MIA Gibbs Sampling
We now provide a theoretical result to show that it is possible to perform the joint sampling and converge to the true target stationary distribution $\Pr(\mathbf{M}|o_{\theta})$, as well as the almost sure convergence of any metric of success of the attack.
@@
as $T\to\infty$. This establishes that the Gibbs sampling procedure for joint MIA converges to the correct joint membership distribution, and empirical averages computed from the samples converge almost surely to their expected values under the target distribution. ∎
@@ B-A Dataset Description
MNIST. MNIST [27] is a benchmark dataset for evaluating handwritten digit classification algorithms. It contains 60,000 grayscale images of 28×28 pixels for training and 10,000 images for testing. The dataset consists of 10 classes, representing digits from 0 to 9.
Fashion-MNIST. Fashion-MNIST (FMNIST) [28] is a dataset designed as a more challenging replacement for MNIST, containing 60,000 grayscale images of size 28×28 pixels for training and 10,000 images for testing. It consists of 10 classes representing different fashion items, such as t-shirts, trousers, and sneakers.
CIFAR10. CIFAR-10 [29] is a benchmark dataset for general image classification tasks, containing 60,000 color images of size 32×32 pixels, equally distributed across 10 classes, including animals like cats, dogs, and birds, as well as vehicles like airplanes and trucks.
CIFAR-100. CIFAR-100 [29] is a dataset similar to CIFAR-10 but with a greater level of complexity, as it contains 100 classes instead of 10. It also includes 60,000 color images of size 32×32 pixels, with each class containing 600 images.
Purchase. Purchase [1] is a dataset of shopping records with 197,324 samples of 600 dimensions. Following previous works [1, 13, 9], we cluster these data into 100 classes to train a Multi-Layer Perceptron (MLP) classifier.
Texas. This dataset is designed to predict a patient’s primary medical procedure using 6,170 binary features. Following [1], the data comprises records from 67,330 patients, with the 100 most frequent procedures used as classification labels.
@@ B-B Accuracy of the Trained Target Model
We report the training and validation accuracies of the target model for both adaptive and non-adaptive settings in Table XIV and Table XV, respectively.
@@ C-A Experiments of $\mathsf{CMIA}$ on Other Models
Here, we present the experimental results on the remaining three model architectures in the adaptive setting. Specifically, Table XVI shows the results for VGG16, Table XVII presents the results for DenseNet121, and Table XVIII displays the results for MobileNetV2. We also present the ROC curves to better demonstrate the improvement achieved by $\mathsf{CMIA}$. Specifically, we use LiRA as the base attack and demonstrate the improvement across four image datasets in Figure 9.
@@ C-B Experiments of $\mathsf{PMIA}$ on Other Models
We present the experimental results on the remaining three model architectures in the non-adaptive setting. Specifically, Table XIX shows the results for VGG16, Table XX presents the results for DenseNet121, and Table XXI displays the results for MobileNetV2. We also present the ROC curves of non-adaptive attacks on four datasets, as shown in Figure 10.

### v3

@@ Abstract
publicationid: pubid: Network and Distributed System Security (NDSS) Symposium 2026 23 - 27 February 2026 , San Diego, CA, USA ISBN 979-8-9919276-8-0 https://dx.doi.org/10.14722/ndss.2026.230661 www.ndss-symposium.org
@@ V-A Experimental Setup
  • LiRA [9] exploits behavioral discrepancies by modeling its loss distribution as a Gaussian estimate and uses a likelihood ratio test to compute the membership score.
@@ V-E Additional Investigations
Attack on Non-image Datasets. We also evaluate the attack performance on two non-image datasets to demonstrate the generality of the proposed attacks. Specifically, we train multilayer perceptrons (MLPs) on the Purchase and Texas datasets [1]. For $\mathsf{CMIA}$, we use LiRA as the base model and apply the cascading framework. For $\mathsf{PMIA}$, we employ the Wasserstein distance to select the top 10 proxy models for each query instance. The attack performance under both adaptive and non-adaptive settings is presented in Table VII. As shown, both $\mathsf{CMIA}$ and $\mathsf{PMIA}$ consistently outperform existing attacks, particularly in TPR at low FPR. It is worth mentioning that the performance of all evaluated MIAs is notably lower on these non-image datasets compared to image datasets such as CIFAR-100, despite both having 100 classes. We attribute this to the smaller generalization gap in models trained on non-image datasets. For instance, as shown in Table XV, the Purchase model achieves training and validation accuracies of 0.96 and 0.80, respectively. In contrast, the CIFAR-100 model shows a much larger gap, with a training accuracy of 0.92 and a validation accuracy of only 0.66. Since MIAs typically exploit overfitting [9], they are less effective against better-generalized models, as observed in the non-image datasets.
@@
Attack Against DP-SGD. Machine learning with differential privacy [43] is an effective defense mechanism against privacy attacks, including MIAs. We follow [9, 8] and assess the effectiveness of the DP-SGD against our attack. Specifically, we fix the clipping norm $C$ as 10 and test $\mathsf{CMIA}$ and $\mathsf{PMIA}$ on a ResNet50 model trained on CIFAR-10. As shown in Table X, even when only the gradient norm is applied without adding noise, the model’s accuracy and the effectiveness of our attack are significantly reduced. We observe that as the noise level increases, the improvement of $\mathsf{CMIA}$ compared to the base attack diminishes. This is because $\mathsf{CMIA}$ relies on identifying highly probable samples to perform cascading attacks on remaining instances. When selecting reliable anchors becomes difficult, the benefit of the cascading framework decreases. We also evaluate the impact of DP-SGD in the non-adaptive setting in Table XI and observe similar patterns: as the privacy level increases, attack performance drops for all methods, and the gap between attacks narrows. Nevertheless, $\mathsf{PMIA}$ still outperforms existing MIAs in the low false-positive regime.
@@ VI Related Work
Non-Adaptive MIAs. Some studies [14, 12] argue that adaptive attacks are inefficient due to the need to train shadow models for every batch of queries. A series of non-adaptive MIAs have been proposed to relax this assumption by training shadow models (or forgoing shadow training entirely) before accessing the query set. Several techniques have been introduced for non-adaptive MIAs, such as score functions [15, 39], difficulty calibration [24, 37, 8], loss trajectory [13, 56], quantile regression [12], and hypothesis testing [38, 14]. However, these attacks deviate from the membership posterior odds ratio test, leading to suboptimal performance.
Recent research [9, 57] has emphasized evaluating attacks by calculating their true positive rate at a significantly low false positive rate, and most existing MIAs perform poorly under this evaluation paradigm. In Section IV-A we provide theoretical insights into the rationale behind these evaluation metrics. While some literature [58] explores instance vulnerabilities to MIAs, it does not propose an attack that leverages them. In contrast, we theoretically analyze membership dependencies and introduce a framework that exploits these vulnerabilities and dependencies to mount a more powerful attack. The dependencies we investigate differ from group-level MIAs [59, 60], which consider the inference target as a set of instances. Instead, our focus lies in exploring the membership dependencies between inference targets.
@@ VII Discussion & Conclusion
In this paper, we provide a new formulation of the MIA game and categorize existing MIAs into two categories: adaptive and non-adaptive. Guided by theoretical analyses of joint membership estimation and the posterior odds test, we propose two attacks (i.e., $\mathsf{CMIA}$ and $\mathsf{PMIA}$), one for each setting. Extensive experiments demonstrate the efficacy of the proposed attacks. Our work has several limitations. First, the conditional shadow training strategy of $\mathsf{CMIA}$ limits its applicability for attacks that do not rely on shadow training. Second, our evaluations mainly focus on classification models, leaving the performance on generative models [61, 62, 63] unexamined. Nevertheless, our work provides both theoretical insights and practical approaches, offering new directions for MIAs.
@@
This paper focuses on membership inference attacks using publicly available datasets that do not contain personally identifiable information. Although our proposed attacks demonstrate improved attack performance, we show that established defenses like DP-SGD remain an effective mitigation strategy. We encourage the research community to use our findings in a manner that promotes ethical research and enhances privacy protections for users and data subjects.
Acknowledgment
We thank the anonymous reviewers for their constructive comments. This work was funded in part by the National Science Foundation (NSF) under awards CNS-2207204, CAREER IIS-1943364, CNS-2212160, and IIS-2229876, as well as by Cisco. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the sponsors.
@@ A-A Convergence of Joint MIA Gibbs Sampling
as $T\to\infty$. This establishes that the Gibbs sampling for joint MIA converges to the joint membership distribution, and empirical averages computed from the samples converge almost surely to expected values under the target distribution. ∎
@@ B-A Dataset Description
MNIST [27] contains 60,000 grayscale images for training and 10,000 images for testing. The dataset consists of 10 classes, representing digits from 0 to 9.
Fashion-MNIST [28] contains 60,000 grayscale images of size 28×28 pixels for training and 10,000 images for testing. It consists of 10 classes representing different fashion items.
CIFAR10 [29] is a benchmark dataset for general image classification tasks, containing 60,000 color images, equally distributed across 10 classes.
CIFAR-100 [29] is a dataset similar to CIFAR-10 but with a greater level of complexity, as it contains 100 classes.
Purchase [1] is a dataset of shopping records with 197,324 samples of 600 dimensions.
Texas [1] comprises records from 67,330 patients, with the 100 most frequent procedures used as classification labels.
@@ B-B Accuracy of the Trained Target Model
The training and validation accuracies of the target model are in Table XIV and Table XV, respectively.
Appendix C Additional Experimental Results
@@ C-A Experiments of $\mathsf{CMIA}$ on Other Models
We present the results on the remaining three model architectures in the adaptive setting. Specifically, Table XVITable XVII and Table XVIII displays the results for VGG16, DenseNet121, and MobileNetV2. We also present the ROC curves of $\mathsf{CMIA}$ using LiRA as the base attack in Figure 9.
@@ C-B Experiments of $\mathsf{PMIA}$ on Other Models
We present the results on the remaining three model architectures in the non-adaptive setting. Specifically, Table XIXTable $X$ and Table XXI displays the results for VGG16, DenseNet121, and MobileNetV2. The ROC curves of non-adaptive attacks are shown in Figure 10.